OpenStudy (anonymous):

How many irrational numbers are there between 1 and 6 ? F. 1 G. 3 H. 4 J. 10 K. Infinitely many

Directrix (directrix):

There are infinitely many non-imaginary numbers between 1 and 6 or between any two numbers for that matter. To see how this works, pick any two of those infinitely many numbers between 1 and 6, multiply them, and then take the square root provided the radicand is not a perfect square. The result will be an irrational number. Example: 7/4 and 9/4 7/4 * 9/4 = 63 / 16 The square root of 63/16 = ( 3* sqrt(7) all over 4 ). This is an irrational number. You can generate as many of these as you like. @tosin Therefore, K. Infinitely many